# Dynamic Memory Compression: Retrofitting LLMs for Accelerated Inference

### v1

@@ 5.3 Per-Head Learned Compression Rates
Since the training loss does not enforce any compression schema a priori, as it just requires to match a certain global CR, we can investigate what schema the model discovers in practice. In Figure 6, we report the CR for each layer and head for different scales (7B, 13B, and 70B) and CRs (2$\times$ and 4$\times$). From all schema, it emerges how compressing deeper layers ($>$ 16 for 7B, $>$ 22 for 13B, $>$ 44 for 70B) is universally the most popular strategy. However, the very final layers are compressed to a somewhat reduced degree. Fascinatingly, 4×4$\times$4 × DMC achieves extremely high CRs for several heads also in the few layers after the very first one. This is counter-productive as token representations are not contextualized yet, which makes taking the correct decision (whether to append or accumulate) hard.
@@
This same pattern (in terms of relative preference for compressing the certain ranges of layers) also corresponds to how the distribution of CRs across layers changes throughout training steps, as we keep annealing the auxiliary loss $\ell_{\text{CR}}$ towards the target CR. Figure 7 in Appendix C illustrates how CR increases first in deeper layers, then in some non-contiguous intermediate ranges.
@@ 6 Related Work
Grouped Query Attention (GQA; Ainslie et al., 2023) represents the most widespread strategy, evolving from Multi Query Attention (MQA; Shazeer, 2019). GQA reduces the number of KV heads by allocating shared KV representations across subsets of query heads. Prior efforts in token merging (Zhang et al., 2018) condensed the entire past context into a single token, while (Liu et al., 2018; Rae et al., 2019) employed strided convolution and mean pooling kernels to reduce the number of KV tokens. Sliding window attention techniques (Beltagy et al., 2020; Child et al., 2019) restrict attention to a maximum of $w$ preceding tokens. Though effective in limiting KV cache, such methods perform fixed-size compression, unlike the presented dynamic DMC, which adapts the compression schema based on the input sequence. This yields superior results, as we prove in an ablation in Appendix E.
@@
Previous learnable compression methods (Anagnostidis et al., 2023, inter alia) decide which tokens to drop from the KV cache. DMC takes a different approach as instead of dropping tokens it merges them. Hence, it preserves cached information more faithfully. Moreover, the DMC compression mechanism has constant complexity with respect to the context length, while the one proposed by Anagnostidis et al. (2023) is linear. Mu et al. (2023) instead compresses prompts through costly generation which limits their inference benefits. Moreover, this method is only applicable to compressing the model input while DMC compresses on-the-fly the entire sequence, including both the model input and the generated output.
@@
Non-learnable cache eviction strategies (Zhang et al., 2023; Sheng et al., 2023; Liu et al., 2023; Wang et al., 2020; Ge et al., 2023) utilize attention scores or token properties to filter tokens in the KV cache. These approaches, while bypassing additional training, rely on heuristics and lack the ability to learn the compression mechanisms. In contrast, DMC integrates compression into its learning objective in an end-to-end manner, where compression is synergistic to language generation.
@@
Finally, DMC draws inspiration from Dynamic Token Pooling (Nawrot et al., 2022), which introduces a learnable boundary predictor to merge the representations of groups of tokens in intermediate layers. DMC improves upon this idea by applying it to the KV cache of and introducing a continuous relaxation of the pooling decision during training. Moreover, it enables retrofitting pre-trained LLMs with minimal extra steps rather than training language models from scratch.
7 Conclusions and Future Work
@@
We proposed Dynamic Memory Compression, a method to reduce the length of the KV cache in Transformers, which enhances the memory efficiency and speed of LLMs at inference time. For every new token, DMC learns end-to-end whether to append its KV representations to the cache or merge them with the top element in the cache. We show how to retrofit LLMs such as Llama 2 at different scales (7B, 13B, and 70B) into efficient DMC versions with a negligible amount of extra data and without extra parameters. DMC LLMs with 2$\times$ and 4$\times$ compression rates (CRs) retain (or even improve upon) the performance of the original LLM. In addition, we demonstrate that, for comparable CRs, DMC has significantly higher continued pre-training performance and sample efficiency than Grouped Query Attention (GQA), a widespread method for KV cache size reduction.
@@
Appendix A Memory-Bound Operations in Transformers
During autoregressive generation with a KV cache, the sequence length during every forward pass is $n=1$. The vast majority of time is spent on calculations for linear layers and multi-head self-attention. For linear layers, the ratio of FLOPs to input bytes improves as the batch size increases. For small batch sizes, the operations are memory bounded on reading the weight matrices from HBM. On the other hand, for MHSA layers during inference the ratio of FLOPs to input bytes does not change and MHSA layers are always memory bounded on current GPUs. The impact of KV cache compression is two-fold as it 1) allows us to decrease the latency of MHSA layers, and 2) allows us to fit larger batch sizes in HBM, resulting in better throughput and better utilization of GPUs during the calculations of linear layers.
Appendix B Replicating the Original Results
To make sure that our implementation is correct, for each downstream task we compare the performance reported in the original Llama 2 paper (Touvron et al., 2023) with those obtained from the Hugging Face Hub checkpoints. Furthermore, we evaluate the impact of using our internal data mixture for up-training, acknowledging that variations in data proportions and preprocessing methodologies can influence model behavior. In particular, we up-train the vanilla pre-trained Llama-2 checkpoint for 200 training steps, amounting to 1B tokens, in accordance with the original Llama-2 training schedule. We compute the average and standard deviation of checkpoints after 50, 100, 150, 200 steps. In our experiments, we replicate the results reported by the Llama 2 paper almost exactly, as shown in Table 2.
@@ Appendix C Analysis of the Compression Schema Learned by DMC
In Figure 7, we illustrate how the CR changes for each layer of the Llama 2 7B model throughout the training from 1$\times$ up to 4$\times$ global CR. Each subplot corresponds to a different global CR which occurs at different stages of the training, going from the smallest (1.17) at the top, to the highest (4.16) at the bottom. There is a clear trend such that, for a smaller global Compression Ratio (i.e. at the beginning of the training), the model emphasizes compression in the later layers. As the global Compression Ratio increases, the model keeps on compressing in final layers but also starts to compress the earlier layers. We hypothesize that the token representations in the initial layers do not contain sufficient information to perform any meaningful grouping. Conversely, token representations in the subsequent layers are more defined and, possibly, after several attention layers, already contain redundant/shared information.
@@
Do DMC models compress sequences with a uniform CR independent from their total length? We find that this is not the case. As show by Figure 8, the CR increases logarithmically as we increase the total sequence length. This holds true across all global CRs (including both 2$\times$ and 4$\times$).
@@
Figure 9: Average value of the decision $\alpha$ for positions (0, 4096) averaged over 128 samples, heads and layers.
Compression Schemata learned by DMC-C
Figure 10: Heatmaps of average compression rates across layers (X-axis) and heads (Y-axis) for DMC-C. Heads are arranged from the highest compression to the lowest top-down for clarity.
@@
More generally higher layers merge longer tokens sequences, in line with Figure 6. For instance, Figure 12 shows the decisions of layer 24 head 2. We leave a more in-depth analysis of compression schemata learned by DMC to future work.
@@
Appendix D Training Ablations
Training Steps per Increase in CR
Another advantage of DMC is its high flexibility. In fact, based on the availability of resources, different regimes can be chosen when annealing the CR to the target during up-training. In Figure 13, we compare a Short and a Long regime for the contrained variant of DMC (DMC-C), which continuously increase the CR by 1 every 3K and 6K steps (12B and 24B tokens), respectively. It is evident how there exists a trade-off between training steps (hence, time) and performance. Additionally, Figure 13 showcases another aspect of the higher flexibility DMC affords: it is compatible with arbitrary real-valued CRs, as opposed to integer CRs divisible by 2 as in GQA.
Schedules of Target CR
Additionally, we explore how different schedules for the target CR impact the model performance. In the standard setup, this is annealed from 1 to the target CR throughout the duration of training. Here, we compare it with a setup where the CR used in the auxiliary loss for compression is set to the target from the start (DMC-immediate). We show the results in Figure 14. As expected, DMC-immediate has a perplexity spike at the beginning when the model quickly increases the CR due to the auxiliary loss. While perplexity is recovered during training, even to a lower point than DMC with annealing, downstream accuracy on MMLU benchmark is degraded across the board. This showcases why avoiding perplexity spikes is fundamental to successfully retrofit an LLM.
Appendix E DMC Ablations
Figure 15: Validation perplexity and test MMLU accuracy vs compression rate for different types of compression priors, in-layer relaxations, and importance scores. All models follow the short training regime and are trained up to 2$\times$ CR.
Fixed vs Learned Memory Compression
We assessed the importance of dynamically learning compression decisions in DMC by comparing it with Fixed Memory Pooling, which reduces the overall number of tokens in memory by deterministically averaging every $n$ tokens, where $n$ in this case is identical to the compression rate. The results, shown in Figure 14, demonstrate that the dynamic component of DMC is crucial to achieve lower perplexity as well as higher downstream accuracy.
Global vs Local (Layer-wise) Compression Prior
In-layer Relaxation
@@ Appendix E DMC Ablations
Table 3: MMLU accuracy (Acc.), Commonsense Question Answering (CS-QA) Exact Match (EM), and Human-Eval Pass@1 for several scales (7B, 13B, and 70B) and compression rates (CRs; 1×1$\times$1 ×, 2×2$\times$2 ×, and 4×4$\times$4 ×) of Llama 2. Here, we include an extra DMC variant - DMC-C, which does not require custom implementation. (*) The 70B model was trained with GQA which compresses the KV cache 8×8$\times$8 ×.
@@
Appendix F Masking Implementation Details
@@
Appendix G Limitations
This paper is focused on retrofitting existing LLMs into DMC variants. In our preliminary experiments with pre-training LLMs with DMC from scratch we obtained negative results when compared to the training curve of GQA. We speculate that this is due to the mutual dependency of modeling and segmenting data: when token representations are not of sufficient quality, boundary decisions are unreliable. Vice versa, incorrect boundary decisions may lead to poor token representations. This creates a vicious cycle which may be broken by techniques that facilitate convergence, such as an Expectation Maximization-style alternation between modeling and segmenting.

### v2

@@ 5.3 Per-Head Learned Compression Ratios
Since the training loss does not enforce any compression schema a priori, as it just requires matching a certain global CR, we can investigate what schema the model discovers in practice. In Figure 6, we report the CR for each layer and head for different scales (7B, 13B, and 70B) and CRs (2$\times$, 4$\times$, 6$\times$, and 8$\times$). From all schema, it emerges how compressing deeper layers ($>$ 16 for 7B, $>$ 22 for 13B, $>$ 44 for 70B) is universally the most popular strategy. However, the very final layers are compressed to a somewhat reduced degree. Fascinatingly, >4×>4$\times$> 4 × DMC achieves extremely high CRs for several heads also in the first few layers. This is arguably counter-productive as token representations are not contextualized yet, which could make decisions (whether to append or accumulate) sub-optimal.
Figure 6: Heatmaps of average compression ratios across layers (x-axis) and heads (y-axis). Heads are arranged from the highest compression to the lowest top-down for clarity.
@@
This same pattern of relative preference for compressing certain ranges of layers also emerges during training as we push the model towards the target CR with the auxiliary loss $\ell_{\text{CR}}$. Figure 7 in Appendix C illustrates how the effective CR increases first in deeper layers, then in some non-contiguous intermediate ranges.
@@ 6 Related Work
Grouped Query Attention (GQA; Ainslie et al., 2023) represents the most widespread strategy, evolving from Multi Query Attention (MQA; Shazeer, 2019). GQA reduces the number of KV heads by allocating shared KV representations across subsets of query heads. Multi-Head Latent Attention (DeepSeek-AI, 2024) compresses the size of the token KV states through a low-rank projection shared across all query heads. Prior efforts in token merging (Zhang et al., 2018) condensed the entire past context into a single token, while (Liu et al., 2018; Rae et al., 2020) employed strided convolution and mean pooling kernels to reduce the number of KV tokens. Sliding window attention techniques (Beltagy et al., 2020; Child et al., 2019) restrict attention to a maximum of $w$ preceding tokens. Quantization-based approaches (Sheng et al., 2023; Liu et al., 2024, 2023b; Hooper et al., 2024) reduce the KV precision to a smaller number of bits. Though effective in limiting KV cache, such methods perform fixed-size compression, unlike the dynamic DMC presented here, which adapts the compression schema based on the input sequence. This adaptation yields superior results, as we prove in an ablation study in Appendix F.
@@
Previous learnable compression methods (Anagnostidis et al., 2023, inter alia) decide which tokens to discard from the KV cache. DMC takes a different approach; instead of dropping tokens, it merges them, thus preserving cached information more faithfully. Moreover, the DMC compression mechanism has constant complexity relative to the context length, while the method proposed by Anagnostidis et al. (2023) has linear complexity. Mu et al. (2023) instead compresses prompts through a costly generation which limits their inference benefits. Moreover, their method is applicable only to compressing the model input while DMC compresses the entire sequence on the fly, including both the model input and the generated output.
@@
Non-learnable cache eviction strategies (Zhang et al., 2023; Sheng et al., 2023; Liu et al., 2023a; Wang et al., 2020; Ge et al., 2023; Oren et al., 2024) utilize attention scores or token properties to filter tokens in the KV cache. These approaches, while bypassing additional training, rely on heuristics and lack the ability to learn the compression mechanisms. In contrast, DMC integrates compression into its learning objective in an end-to-end manner, where compression is synergistic to language generation.
@@
Finally, DMC draws inspiration from Dynamic Token Pooling (DTP; Nawrot et al., 2023), which introduces a learnable boundary predictor to merge the representations of groups of tokens in intermediate layers. DMC improves upon this idea by applying it to the KV cache and introducing a continuous relaxation of the pooling decision during training. Moreover, it enables retrofitting pre-trained LLMs with minimal extra steps rather than training language models from scratch.
7 Conclusions
@@
We proposed Dynamic Memory Compression, a method to reduce the length of the KV cache in Transformers, which enhances the efficiency of LLMs at inference time. For every new token, DMC learns end-to-end whether to append its key–value representations to the cache or merge them with the top element in the cache. We showed how to retrofit LLMs such as Llama 2 at different scales (7B, 13B, and 70B) into efficient DMC versions with a negligible amount of extra data and without extra parameters. DMC LLMs with up to 4$\times$ compression ratios (CRs) retain (or even improve upon) the performance of the original LLM. Higher CRs (6$\times$ and 8$\times$) instead result in limited performance degradation. For comparable CRs, DMC has significantly higher downstream performance and sample efficiency than Grouped Query Attention (GQA), a widespread method for KV cache size reduction, as well as key–value eviction policies such as H2O and TOVA. In practice, we find that DMC translates to up to 7$\times$ increase in throughput on an H100 GPU.
Acknowledgements
The authors would like to thank Mostofa Patwary for sharing the data blend we used to retrofit the models, Szymon Migacz for his assistance with the computing infrastructure, as well as Przemysław Strzelczyk, Daniel Korzekwa, and Bryan Catanzaro for helpful discussions and support in releasing this paper. This work was supported in part by the UKRI Centre for Doctoral Training in Natural Language Processing, funded by the UKRI (grant EP/S022481/1) and the University of Edinburgh, School of Informatics and School of Philosophy, Psychology & Language Sciences.
@@ Acknowledgements
Appendix A Replicating the Original Results
To make sure that our implementation is correct, for each downstream task, we compare the performance reported in the original Llama 2 paper (Touvron et al., 2023) with those obtained from the Hugging Face Hub checkpoints. Furthermore, we evaluate the impact of using our internal data mixture for up-training, acknowledging that variations in data proportions and preprocessing methodologies can influence model behavior. In particular, we up-train the vanilla pre-trained Llama 2 checkpoint for 200 training steps, amounting to 1B tokens, in accordance with the original Llama 2 training schedule. We compute the average and standard deviation of checkpoints after 50, 100, 150, and 200 steps. In our experiments, we replicate the results reported by the Llama 2 paper almost exactly, as shown in Table 2. Furthermore, we observe that retrofitting the Llama 2 checkpoint on our data mixture has little effect on the model’s performance on the downstream benchmarks.
@@ Appendix A Replicating the Original Results
Appendix B Retrofitting Data
The retrofitting corpus comprised sections of The Pile, including BookCorpus2, Books3, Pile-CC, Gutenberg (PG-19), NIH ExPorter, OpenWebText2, Stack Exchange, and Wikipedia (en). Additional datasets included in the corpus were ArXiv, Pushshift Reddit, mC4 (multilingual C4), Common Crawl (CC) dumps from 2017-2023, CC-News, PubMed Central, BigScience Workshop datasets, and The Stack dataset (BigCode project).
@@ Appendix C Analysis of the Compression Schema Learned by DMC
In Figure 7, we illustrate how the CR changes for each layer of the Llama 2 7B model throughout the training from 1$\times$ up to 4$\times$ global CR. Each subplot corresponds to a different global CR which occurs at different stages of the training, going from the smallest (1.17) at the top to the highest (4.16) at the bottom. There is a clear trend such that, for a smaller global Compression Ratio (i.e. at the beginning of the training), the model emphasizes compression in the later layers. As the global Compression Ratio increases, the model keeps on compressing in the final layers but also starts to compress the earlier layers. We hypothesize that the token representations in the initial layers do not contain sufficient information to perform any meaningful grouping. Conversely, token representations in the subsequent layers are more defined and, possibly, after several attention layers, already contain redundant/shared information.
@@
Do DMC models compress sequences with a uniform CR independent from their total length? We find that this is not the case. As shown by Figure 8, the CR increases logarithmically as we increase the total sequence length. This holds true across all global CRs (including both 2$\times$ and 4$\times$).
@@
Figure 9: Average value of the decision $\alpha$ for positions (0, 4096) averaged over 128 samples, heads, and layers.
@@
More generally higher layers merge longer token sequences, in line with Figure 6. For instance, Figure 11 shows the decisions of layer 24 head 2. We leave a more in-depth analysis of compression schemata learned by DMC to future work.
@@
Appendix D Similar Per-layer Compression Rates
In order to alleviate the problem of efficiently storing and retrieving the KV cache of variable-length sequences, we also study a Constrained variant of our algorithm (DMC-C) in which we force all heads in a given layer to maintain a similar compression ratio. A similar compression ratio allows us to store key and value sequences naïvely as padded tensors with minimal padding. To this end, we add an extra loss term to Equation 11
$\ell_{\text{H}}=\sum_{l=1}^{n_{l}}\sum_{h=1}^{n_{h}}\sum_{t=1}^{n}\left|\alpha_{lht}-\sum_{h=1}^{n_{h}}\alpha_{lht}\right|.$ (14)
@@ Appendix D Similar Per-layer Compression Rates
Table 3: MMLU accuracy, Commonsense Question Answering (CS-QA) accuracy averaged across 6 tasks, and HumanEval Pass@1 for several scales (7B, 13B, and 70B) and compression ratios (CRs; 1×1$\times$1 ×, 2×2$\times$2 ×, and 4×4$\times$4 ×) of Llama 2. Here, we include an extra DMC variant - DMC-C, which does not require custom implementation. (*) The 70B model was trained with GQA which compresses the KV cache 8×8$\times$8 ×.
Table 3 compares DMC with its Constrained variant DMC-C. In general, while remaining superior to GQA, DMC-C displays a significant degradation in several configurations, most notably 7B 4$\times$ where it records a drop of -6.4 in MMLU compared to the ceiling. On the other hand, DMC recovers all performance loss in DMC-C. When combined with custom attention implementations that do not require excessive padding, standard DMC should therefore be vastly preferred, as it retains the original LLM performance while fully reaping the advantages in memory efficiency.
Compression Schemata learned by DMC-C
Figure 12: Heatmaps of average compression ratios across layers (x-axis) and heads (y-axis) for DMC-C. Heads are arranged from the highest compression to the lowest top-down for clarity.
Appendix E Training Ablations
Training Steps per Increase in CR
Another advantage of DMC is its high flexibility. In fact, based on the availability of resources, different regimes can be chosen when annealing the CR to the target during up-training. In Figure 13, we compare a Short and a Long regime for the constrained variant of DMC (DMC-C), which continuously increases the CR by 1 every 3K and 6K steps (12B and 24B tokens), respectively. Evidently, there exists a trade-off between training steps (hence, time) and performance. Additionally, Figure 13 showcases another aspect of the higher flexibility DMC affords: it is compatible with arbitrary real-valued CRs, as opposed to integer CRs divisible by 2 as in GQA.
Schedules of Target CR
Additionally, we explore how different schedules for the target CR impact the model performance. In the standard setup, this is linearly annealed from 1 to the target CR throughout the duration of training. Here, we compare it with a setup where the CR used in the auxiliary loss for compression is set to the target from the start (DMC-immediate). We show the results in Figure 14. As expected, DMC-immediate has a perplexity spike at the beginning when the model quickly increases the CR due to the auxiliary loss. While perplexity is recovered during training, even to a lower point than DMC with annealing, downstream accuracy on the MMLU benchmark is degraded across the board. This showcases why avoiding perplexity spikes is fundamental to successfully retrofitting an LLM.
Appendix F DMC Ablations
Figure 15: Validation perplexity and test MMLU accuracy vs compression ratio for different types of compression priors, in-layer relaxations, and importance scores. All models follow the short training regime and are trained up to 2$\times$ CR.
Fixed vs Learned Memory Compression
We assessed the importance of dynamically learning compression decisions in DMC by comparing it with Fixed Memory Pooling, which reduces the overall number of tokens in memory by deterministically averaging every $n$ tokens, where $n$ in this case is identical to the compression ratio. The results, shown in Figure 14, demonstrate that the dynamic component of DMC is crucial to achieving lower perplexity as well as higher downstream accuracy.
Global vs Local (Layer-wise) Compression Prior
In-layer Relaxation
@@ Appendix F DMC Ablations
Appendix G Masking Implementation Details
@@
Appendix H Limitations
This paper is focused on retrofitting existing LLMs into DMC variants. In our preliminary experiments with pre-training LLMs with DMC from scratch, we obtained negative results when compared to the training curve of GQA. We speculate that this is due to the mutual dependency of modeling and segmenting data: when token representations are not of sufficient quality, boundary decisions are unreliable. Vice versa, incorrect boundary decisions may lead to poor token representations. This creates a vicious cycle that may be broken by techniques that facilitate convergence, such as an Expectation Maximization-style alternation between modeling and segmenting.